Group Selection Page HELP Controls Group ITDM Vacuum WP78 FXE SPB SQS SXP XO Migrated2Zulip Archived2GitlabPages Archived2PDF not2MigrateOrArchive MigratedByGroup
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB_mig
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 278     Entry time: 28 Jul 2023, 15:37
Author: Harald Sinn 
Group: PRC 
Subject: Status 

Good news: The 'jitter'-tests was sucessful and BKR could identify the jitter source related to a wrong software setting around the AH1 module of injector.

This was now fixed and the jitter has disappeared smiley . It is likely that the tuning was compromised by this setting from the beginning of this run (July11).

On Monday morning, a fresh tuning round will be started with hopefully improved SASE performance. Until then photon delivery with what we have.  

ELOG V3.1.4-7c3fd00